G3tH3Pxmpthat5Amx5t:

Which expression is equivalent to (f + g)(4)?

dude:

We'll need to know what f(x) and g(x) are

G3tH3Pxmpthat5Amx5t:

Is it one of these?

G3tH3Pxmpthat5Amx5t:

f(4) + g(4) f(x) + g(4) f(4 + g(4)) 4(f(x) + g(x))

dude:

Ah I see I mean any guesses? (f+g) is adding both equations where they both have 4 = x (you'd want to split them and have both include 4)

G3tH3Pxmpthat5Amx5t:

o it would be the 3 1?

dude:

No no, you want 4 in g(x) and f(x) Ex. \((f*g)(7) = f(7) * g(7)\)